OpenStudy (anonymous):

How deep can a snorkel diver dive if the maximum gauge pressure on the thorax cannot exceed 11.6 kPa? Hint: consider sea water of density 1.061 g/cm3.

OpenStudy (anonymous):

using the equation p = height * density* g 11600 = 1061*10* h h = 11600/10610 =1.09 m